Week of 190408

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Alberto (Monitoring), Ben (Computing), Borja (Chair, Monitoring), Ivan (ATLAS), Mark (LHCb), Remi (Storage), Vincent (Security)
  • remote: Andrew (Nikhef), Di (TRIUMF), John (RAL), Marcelo (CNAF), Sang-Un (KISTI), Dave (FNAL).

Experiments round table:

  • CMS reports ( raw view) -
    • Due to CMS week ongoing, we likely have nobody joining the call today
    • Good resource utilization ~230k cores
      • Usual mix 25% Analysis, 75% Production
    • No major issues

  • ALICE -
    • Apologies: probably no ALICE expert will attend the meeting
    • High activity in the past week
    • No major issues, at least until Monday morning

  • LHCb reports ( raw view) -
    • Activity
      • User jobs, MC productions, staging and some reprocessing starting this week.
    • Issues
      • RAL:
        • A restart of docker killed a number of jobs last week. RAL investigating the course ( GGUS:140589)
        • A disk server was in a bad state that caused timeouts on opening some files (GGUS:140599)

Sites / Services round table:

  • ASGC: NC
  • BNL: NTR
  • CNAF: NTR
  • EGI: NC
  • FNAL: NTR
  • IN2P3: NC
  • JINR: NTR
  • KISTI: Scheduled downtime for Monday the 15th from 1:00 AM - 5:00 AM for maintenance of the bridge area network switches. Expects to be transparent but might have some effect on the users.
  • KIT: NTR
  • NDGF: NC
  • NL-T1:
    • Last Wednesday, Surfsara had a problem with a storage array in the tape system. Staging was delayed for a day.
  • NRC-KI: NC
  • OSG: NC
  • PIC: NC
  • RAL:
    • The disk server problem for LHCb continues. We are trying to recover files from the machine (gdss700).
    • The docker restart issue affected all VOs - only LHCb noticed.
  • TRIUMF: NTR

  • CERN computing services:
    • lxplus.cern.ch alias moved to CC7 (SLC6 lxplus6.cern.ch still there) / Default for local HTCondor jobs switched to CC7
    • Continuing to migrate capacity over to CC7 as per OTG:0047300

- Ivan asked if the migration will start converting CEs soon, Ben replied that it's the plan, Ivan wanted to be notified when it started.
- Mark asked about the end of life for SLC6, Ben mentioned that initial plan was to left 10% of the capacity from the 1st of June, but they were asked to be more aggressive in the migration so 1st of June there won't be any more SLC6 batch nodes, but in any case there will be a follow up before the 1st of June on people still using them.
- Ivan mentioned that Lxplus usage is a bit different and whether the migration will follow the same steps. Ben will make sure with them.

  • CERN storage services: NTR
  • CERN databases: NC
  • GGUS: NTR
  • Monitoring:
    • Draft reports for the March availability sent around
  • MW Officer: NC
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r20 < r19 < r18 < r17 < r16 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r20 - 2019-04-09 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback